HW4: Problem 3 Previous Problem List Next (1 point) You can receive one of the following two payment streams: 0 $150 at time 0, $300 at at time n and $350 at time 2n () $800 at time 10. At an annual effective interest rate of i the present values of the two streams are equal. Given v 0.738 determine i Preview My Answers Submit Answers You have attempted this problem 0 times. You have unlimited attempts remaining Email instructor
Answer #1

GIven, present value of two streams are equal

Hence, 150+300/(1+i)^n+350/(1+i)^(2*n)=800/(1+i)^10

Given, v^n=0.738

or, 1/(1+i)^n=0.738

Substituting the value of 1/(1+i)^n in the above equation, we get

150+300*0.738+350*(0.738^2)=800/(1+i)^10

=>i=3.59371069430628%